# DRAFT- Rotation Curve Fitting Model

### v2

@@ Abstract
One key piece of evidence for dark matter is the flat rotation curve problem: the disagreement between measured galactic rotation curves and their luminous mass. A novel solution to this problem is presented here. A model of relativistic frame effects on Doppler shifts due to the slightly curved frames of an emitting galaxy with respect to the Milky Way frame is derived. This model predicts observed Doppler shifted spectra based only on the observed luminous matter profile and one free model parameter. The rotation curve fitting model presented is tested on the SPARC sample of 175 galactic rotation profiles using accurate photometry measurements which has previously been fitted by dark matter models, MOND and the Radial Acceleration Relation. For the SPARC sample of 175 galaxies, the new model fits 172 galaxies with an averaged reduced , the isothermal dark matter model fits 165 galaxies with an average , and the Radial Acceleration Relation fits all 175 galaxies with an average . The model presented here has fewer free parameters than either dark matter models or MOND, and does not modify classical physics. Implications of this model are discussed.
1 Introduction
@@
The flat-rotation curve problem is the divergence of two rotation velocities about the center of a spiral galaxy, inferred from different observations of light (Rubin et al., 1980; Bosma, 1981; van Albada et al., 1985): photometry and Doppler shifted spectra. Photometry gives estimates of the luminous mass, which when interpreted classically by the Poisson equation gives the expected, Keplerian orbital velocities which decline beyond the stars. Doppler shifted spectra, however, gives the “flat-rotation curve” velocities, which remain essentially constant far past the stars. The divergence of these two velocities is primary evidence for dark matter theories (Fig. 1).
@@
Dark matter is hypothesized to be massive particles which are electromagnetically neutral, and have a very low interaction probability with baryonic matter, but otherwise obey classical gravity. Though direct detection experiments continue, in the absence of a definitive observation of such particles (Cebrián, 2022), the paradoxes of dark matter theory remain interesting. S. McGaugh notes one such paradox when he asks “Why is the luminous tail wagging the dark matter dog, if dark matter dominates dynamics?” (McGaugh, 1999). This references the curious fact that knowledge of the stellar disk completely determines the spherical dark matter halo, even though dark matter halos are postulated to be ten times the mass of the luminous matter on average (McGaugh, 2004).
@@
Another interesting paradox is the Universal Rotation Curve (URC), in which a spectrum of 1,100 rotation curves (RCs) inflect about the presumed rotation curve of the Milky Way (Persic et al., 1996; Rubin et al., 1978; Salucci et al., 2007). In the URC spectrum, galaxy RCs are separated with respect to the luminous mass of the Milky Way, where the RCs of galaxies larger than the Milky Way inflect downwards “flat”, and RCs of galaxies smaller than the Milky Way inflect upwards towards “flat”. The Milky Way has been postulated to be at the inflection point where truly flat RCs lie, though a definitive rotation curve past our position at approximately kpc has been difficult to constrain until the current era of observations from Gaia DR3 (Jiao, Yongjun et al., 2023). We interpret this positioning of the Milky Way as evidence for frame dependent effects in this problem.
@@
Dark matter theory accounts for the curious URC phenomenology with the phrasing that galaxies smaller than the Milky Way are “dark matter dominated”, and those larger than the Milky Way require only “minimal dark matter halos”. Dark matter particles are designed to obey classical gravity, and so this logical inconsistency means that dark matter models require fine-tuning to account for the URC phenomenology, reducing the predictivity of the models (McGaugh, 2021). In classical gravity, mass accretion rates are directly proportional to the initial mass function (Maschberger et al., 2014). Interpreting the URC, instead, from the paradigm of relative frame effects removes this conflict, explains why galaxy RCs are binned as they are with respect to the Milky Way in the URC, removes extraneous free parameters and does not modify classical gravity theory.
@@
To construct a paradigm of relative frame effects, we look to the leading alternative to dark matter, Modified Newtonian Dynamics (MOND) (Milgrom, 1983). MOND successfully fits a diverse distribution of galaxy RC, by modifying the standard gravitational acceleration scale on the length scales of galaxies (McGaugh, 2014). In MOND the luminous mass is the only mass, but excesses in Doppler shifted spectra are still interpreted as physical rotation velocities, as in dark matter theories. It is remarkable in that MOND’s free parameter is essentially constant across a wide range of galaxy morphologies (Lelli et al., 2016), though there are notable exceptions where MOND fits of some well-studied galaxies indicate distances which are in conflict with independent measurements from standard candles. Previous attempts to extend MOND to the relativistic regime have given rise to new theories of gravity (Bekenstein, 2004; Cooperstock & Tieu, 2007; Famaey & McGaugh, 2012). However, we suggest a more conservative approach, in which the frame-dependent picture presented here is viewed as a relativistic extension of MOND. The paradigm of relative frame effects does not modify classical physics, and can be seen as subsuming the role of the MONDian changing acceleration scale within the picture of changing relative curvature. This is also why we suggest that MOND is limited, because the single free parameter for a new acceleration scale can not adapt as a function of radius to the changing relative curvature of the Milky Way with respect to the galaxy being observed.
@@
The new rotation curve fitting formula presented here fits the same galaxies as MOND and dark matter, at reported distances, with fewer free parameters, and no modification of classical gravitational physics. The only inputs to the RCFM fits are estimates of the luminous mass of the galaxy being observed and that of the Milky Way. Luminous mass models are those reported in the SPARC database (Lelli et al., 2016), with the addition of a few external models for the sake of comparison in figures. Recent observations of the Milky Way from the Gaia DR3, indicate a Keplerian decline in rotation velocities to kpc, consistent with the model paradigm presented here Jiao, Yongjun et al. (2023).
@@
The new rotation curve fitting model (RCFM) is derived in Section 2. Previous consideration of relative gravitational frame effects in the flat-rotation curve problem used Galilean differences of redshifts at the large limit of the observations (Misner et al., 1973), instead of the Lorentz-type maps used here. In Section 3 we describe the SPARC database of 175 well known galaxy rotation curves and associated luminous mass models, and the Milky Way baryon models used in this work. In Section 4 we compare RCFM fit results to those from dark matter and the Radial Acceleration Relation (RAR) models for the SPARC galaxies. The RAR is an empirical relation (McGaugh, 2014; McGaugh et al., 2016; Lelli et al., 2016; Li et al., 2018), postulated to be the phenomenological basis of MOND (Dutton et al., 2019). In this section we also explore a correlation between the RCFM free parameter and a ratio of photometric observables (luminosity and half light radius of the galaxy). In Section 5 we conclude with discussion of upcoming observations which can further constrain or falsify this model.
@@ 2.1 Two observations of light
The observables in the flat-rotation curve problem are the Doppler shifted spectra and photometry. Doppler shifted spectra give RC velocities by Lorentz boost
for the characteristic lab frequency and the observed shifted frequency. The vacuum light speed is . We emphasize, this form of the Lorentz boost is applicable to a flat, inertial spacetime.
The second observable, photometry, gives a measure of total light which is then interpreted as mass from population synthesis models (PSMs) which produce mass-to-light ratios and hence models of the baryon distribution in a given galaxy. PSM rely upon a complex suite of assumptions regarding galaxy evolution, metallicities and initial mass functions (Bell & de Jong, 2001; Schombert et al., 2018), and are under-constrained due to the dark matter problem (Conroy et al., 2009; Dutton et al., 2005), producing an error budget of (Lelli et al., 2016).
@@
for the field point, the reference point, and the acceleration due to the gravitational field. The classical boundary condition is that the potential goes to zero at , hence the integration effectively starts from large into the small limit of the data. This potential then solves the Poisson equation
@@
The baryonic components are summed in quadrature as stellar disk, stellar bulge, and a gas halo
where and are the stellar mass-to-light ratios, which are free parameters in rotation curve fits. Gas fractions are calculated from a different observational technique as given in Casertano (1983), and do not require mass-to-light ratios. In this way, Eq. 4 can clearly be seen to be a sum of gradients, as is the dark matter RC formula in the next section.
@@ 2.2 Dark matter rotation curve fitting formula
The dark matter rotation curve (RC) formula is
where terms in are the Keplerian velocity predictions in Eq. 5, are the velocities attributed to dark matter, and terms in are the model prediction which are fitted to the RC velocities in Eq. 1 from shifted spectra. All velocities are assumed to be those of test particles in circular orbits, in the plane of the stellar disk, about the rotation axis of the galaxy at .
@@ 2.3 Radial Acceleration Relation
The Radial Acceleration Relation (RAR) is an empirical relation (McGaugh et al., 2016) which has been considered as the phenomenological basis for MOND (Dutton et al., 2019; Desmond et al., 2023). The RAR approaches rotation curve interpretation from the MOND perspective of a changing law of inertia. In what follows, we will use the RAR in comparison to RCFM and dark matter fits, as it has been used to fit the SPARC sample and is a simpler form. In the RAR, the changing law of inertia is stated as
@@ 2.4 New rotation curve fitting model (RCFM)
The physical paradigm for the new rotation curve fitting model (RCFM) is: the luminous mass is the only mass, the velocity due to the baryons is the only velocity, and excesses in Doppler shifted spectra represented by the “flat” rotation velocities are due to relative frame effects from the Milky Way galaxy. For clarity in what follows, we now drop explicit functional dependence on . All terms can be assumed to be evaluated one-to-one in radius between the galaxy being observed and the Milky Way, with the exception of the model’s free parameter which is single valued for each galaxy fitted. Constants like Newton’s and the speed of light will be set to 1.
We construct the heuristic new rotation curve fitting formula by replacing the dark matter contribution to the rotation curve in Eq. 6 with two Lorentz-type maps and which represent the frame-dependent effects due to the Milky Way;
@@
This prescription is an implicit assumption that contributions to Doppler shifted spectra from translation ( in Eq. 5) and from relative curvature () are roughly separable as in Eq. 8 (Jackson, 1999).
@@
In the frame field formalism “when going from one local inertial frame at a given point to another at the same point, the fields transform with respect to a Lorentz transformation” (Jetzer, 2017). Our goal then, is to use Lorentz-type functions to map the target galaxies onto the Milky Way one-to-one in radius, using the weak-field Schwarzschild time coefficients to characterize the clock field which defines the effects on photon frequency. The weak-field timelike Schwarzschild field frames are defined by
Terms in are the Newtonian gravitational potential (Wald, 1984). Classically, the of galaxies are integrated from the large limit of the data with a boundary value of , into a negative maxima at the small limit(Eq. 2). This is an implicit assumption of a globally flat embedding space, which ensures that the potential goes to zero at a distance of infinity from the mass distribution. Recent observations (Pomarède et al., 2020), however, have shown that on the relevant length scales of galaxies and groups of galaxies we have not recovered flat spacetimes. In fact, the external environments of galaxies and flowlines of groups of galaxies are exceedingly complex and diffusely populated with matter, demonstrating structure with sources and sinks. In addition, the value of the vacuum energy remains an outstanding physics problem (Zel’dovich, 1968). In the absence of definite knowledge of the values of at the large limit of individual galaxies, and since all RCFM terms are ratios of terms in for two galaxies, we instead inform the integration of the gravitational potentials from Wolfgang Rindler’s statement that “the center of each galaxy provides a basic local standard of nonacceleration … so then can be treated like a local inertial frame relative to its own center”(Rindler, 2013).
From Rindler’s insight, we compare galaxies as inertial frames by comparing them from their centers. In practice this means that we integrate galaxy potentials from a value of zero at the center where , as we know curvatures are all the same at that point. Computationally, this produces positive definite gravitational potentials, summed from at out to a positive, constant value at the large limit of the data. These potentials are then subtracted from in the clock term . Gravitational potentials calculated in this way still obey Poisson’s equation Eq. 3 and the central force law Eq. 4, but allow for a non-flat embedding space with no loss of generality.
The curved 2-frame mapping is
with a rapidity defined by
for representing the target galaxy and the Milky Way. The second transformation is
for the rapidity angle defined by
for the convolution of the curved 2-frame (Eq. 13) with the flat 2-frame
Terms in are the Keplerian velocities estimated from photometry of the luminous mass. This last step is necessary because observations are made in locally flat frames, where the metric is Minkowski . The Keplerian velocities in Eq. 16 are our best estimate of flatness, as evidenced by the fact that dark matter is not required to reproduce the RC of our Solar System. Previously, Rindler extended the transforms of special relativity to accelerated frames on a flat background (Boyer, 2011); viewed as Rindler’s accelerated coordinates, is timelike and is spacelike (Misner et al., 1973; Wald, 1984; Rindler, 2013).

### v4

# Rotation Curve Fitting Model

@@ Abstract
One key piece of evidence for dark matter is the flat rotation curve problem: the disagreement between measured galactic rotation curves and their luminous mass. A novel solution to this problem is presented here. A model of relativistic frame effects on Doppler shifts due to the slightly curved frames of emitting galaxies with respect to the Milky Way frame is derived. This model predicts observed Doppler shifted spectra based only on the luminous matter estimates and one free model parameter. The rotation curve fitting model presented is tested on the SPARC sample of 175 galactic rotation profiles, also fitted by also fitted by dark matter models, and MOND via the Radial Acceleration Relation. For the SPARC sample, the new rotation curve fitting model gives an averaged reduced for 172 galaxies fitted, the isothermal dark matter model gives for 165 galaxies fitted, and the Radial Acceleration Relation (MOND precursor) gives for 175 galaxies fitted. The model presented here has a free parameter that is highly correlated with a ratio of photometric quantities. Implications of this model are discussed.
keywords:
gravitation < Physical Data and Processes – (cosmology:) dark matter – galaxies: fundamental parameters – galaxies: haloes – galaxies: bulges – galaxies: photometry1 Introduction
@@
The flat-rotation curve problem is the divergence of two rotation velocities about the center of a spiral galaxy, inferred from different observations of light (Rubin et al., 1980; Bosma, 1981; van Albada et al., 1985): photometry and Doppler shifted spectra. Photometry gives estimates of the luminous mass, which when interpreted classically by the Poisson equation gives the expected Keplerian orbital velocities which decline beyond the stars. Doppler shifted spectra, however, give the “flat-rotation curve” velocities, which remain essentially constant far past the stars. The divergence of these two velocities has been primary evidence for dark matter theories (Fig. 1).
@@
Dark matter is hypothesized to be massive particles which are electromagnetically neutral, and have a very low interaction probability with baryonic matter, but otherwise obey classical gravity. Though direct detection experiments continue, in the absence of a definitive observation of such particles (Cebrián, 2022), the phenomenological details of dark matter problems remain interesting. S. McGaugh notes a degeneracy between the luminous and dark matter components in a galaxy, when he asks “Why is the luminous tail wagging the dark matter dog, if dark matter dominates dynamics?” (McGaugh, 1999). This references the curious fact that knowledge of the stellar disk completely determines the spherical dark matter halo, even though dark matter is postulated to dominate dynamics (McGaugh, 2004).
@@
Another interesting trend is the Universal Rotation Curve (URC), in which a spectrum of 1,100 rotation curves (RCs) inflect about the presumed rotation curve of the Milky Way (Rubin et al., 1978; Persic et al., 1996; Salucci et al., 2007). In the URC spectrum, galaxy RCs are separated with respect to the luminous mass of the Milky Way, where the RCs of galaxies larger than the Milky Way inflect downwards towards “flat”, and RCs of galaxies smaller than the Milky Way inflect upwards towards “flat”. The Milky Way has been postulated to be at the inflection point where truly flat RCs lie, though a definitive rotation curve past our position at approximately kpc has been difficult to obtain until the current era of observations from Gaia DR3 (Jiao, Yongjun et al., 2023). We interpret this positioning of the Milky Way as evidence for frame dependent effects in this problem.
@@
Dark matter theory accounts for the curious URC phenomenology with the phrasing that galaxies smaller than the Milky Way are “dark matter dominated”, and those larger than the Milky Way require only “minimal dark matter halos”. Dark matter particles are designed to obey classical gravity, and so this logical inconsistency requires fine-tuning of extra free-parameters, reducing the predictivity of the models (McGaugh, 2021). In classical gravity, mass accretion rates are directly proportional to the initial mass function (Maschberger et al., 2014).
In this paper we interpret the URC phenomenology, instead, as due to relative frame effects from our Milky Way, which immediately explains our apparent position in the middle of a spectrum of 1,100 rotation curves. This interpretation of the problem removes extraneous free parameters, and does not modify classical gravity theory. Modified Newtonian Dynamics (MOND) (Milgrom, 1983), the leading alternative to dark matter, is similar to what we present in so far as MOND states that the luminous mass is the only mass, but differs from our model as it modifies classical gravitational physics to obtain rotation curve velocities. We will instead interpret excess rotation velocities, above the estimates from photometry, as due to an incomplete interpretation of Doppler shifted spectra, not true orbital speeds.
@@
MOND is instructive in this context because it successfully fits a diverse distribution of galaxy rotation curves (RCs), by modifying the standard gravitational acceleration scale on the length scales of galaxies (McGaugh, 2014). Previous attempts to extend MOND into the relativistic regime have given rise to new theories of gravity (Bekenstein, 2004; Cooperstock & Tieu, 2007; Famaey & McGaugh, 2012), but what we propose is simpler, a transition of the concept of a changing acceleration to the relativistic regime of a changing relative curvature. It is remarkable that MOND’s free parameter is essentially constant across a wide range of galaxy morphologies (Lelli et al., 2016), which we interpret as a characterization of the role of the Milky Way in this problem. However, there are some notable exceptions where MOND fits of well-studied galaxies indicate distances which are in conflict with standard candles, and this problem is removed in our formulation of the flat-rotation curve problem.
@@
The new rotation curve fitting model (RCFM) presented here fits the same galaxy data as MOND and dark matter models, at reported distances, but with fewer free parameters and no modification of physics. The RCFM model is a one-parameter fit to the rotation curve data, whose only inputs are estimates of the luminous mass of the galaxy emitting the photons and of galaxy receiving those photons. Recent observations of the Milky Way from the Gaia DR3 (Jiao, Yongjun et al., 2023) indicate a Keplerian decline in rotation velocities from kpc to kpc, consistent with the model paradigm presented here.
@@
The new rotation curve fitting model (RCFM) is derived in Section 2. Historically, gravitational redshifts considered in this problem were obviated by Galilean subtraction of redshifts at the large limit of the data of the emitter galaxy from the receiver galaxy (Misner et al., 1973). We instead employ a relativistic framing, in which field frames are parametrized by the Schwarzschild gravitational redshift terms, and then we Lorentz map between between the emitter and receiver galaxies one-to-one in radius (See Section 2.4). In Section 3 we describe the SPARC database (Lelli et al., 2016) of 175 well known galaxy rotation curves and associated luminous mass models, and the Milky Way baryon models used in this work. In Section 4 we compare RCFM fit results to those from dark matter models and the Radial Acceleration Relation (RAR) (McGaugh, 2014; McGaugh et al., 2016; Lelli et al., 2016; Li et al., 2018) for the SPARC galaxy database (RAR has a simpler functional form and is a possible precursor for MOND based on the same physical paradigm (Dutton et al., 2019)). In this section we also explore a correlation between the RCFM free parameter and a ratio of photometric observables (luminosity and half light radius of the galaxy). In Section 5 we conclude with discussion of upcoming observations which can further constrain or falsify this model.
@@ 2.1 Two observations of light
The observables in the flat-rotation curve problem are the Doppler shifted spectra and photometry.
Doppler shifted spectra is interpreted as RC velocities by the Lorentz boost
for the characteristic lab frequency, the observed shifted frequency, and , for the vacuum light speed. We emphasize this form of the Lorentz boost is applicable to a flat, inertial spacetime.
The second observable, photometry, gives a measure of total light which is then interpreted as mass from population synthesis models (PSMs) which produce mass-to-light ratios and hence models of the baryon distribution in a given galaxy. PSM rely upon a complex suite of assumptions regarding galaxy evolution, metallicities and initial mass functions (Schombert et al., 2018; Bell & de Jong, 2001), and are under-constrained due to the dark matter problem (Conroy et al., 2009; Dutton et al., 2005), producing an error budget of (Lelli et al., 2016).
@@
for the field point, the reference point, and the acceleration due to the gravitational field. The classical boundary condition is that the potential goes to zero at , hence the integration effectively starts from large limit of the data, and integrates in to the small . This potential then solves the Poisson equation
@@
The baryonic components of a galaxy include the stellar disk, stellar bulge, and a gas halo, which are summed in quadrature to reflect a sum of gradients
Mass-to-light ratios and are free parameters in most rotation curve fits due to the under-constrained nature of luminous mass modeling. Gas fractions are calculated from a different observational technique (Casertano, 1983), and do not require mass-to-light ratio modeling.
@@ 2.2 Dark matter rotation curve fitting formula
The dark matter rotation curve (RC) formula is formulated in the same sense as Eq. 5, as a sum of gradients in the potential
for terms in the Keplerian velocity predictions (Eq. 5), the velocities attributed to dark matter, and the model predictions which are fitted to the RC velocities reported from observations of Doppler shifted spectra (Eq. 1). All velocities are assumed to be those of test particles in circular orbits, in the plane of the stellar disk, about the rotation axis of the galaxy at .
@@ 2.3 Radial Acceleration Relation
The Radial Acceleration Relation (RAR)(McGaugh et al., 2016) has been considered as the phenomenological basis for MOND (Dutton et al., 2019; Desmond et al., 2023), as it has a simpler functional form and less scatter than MOND but the same physical paradigm. For this reason, we will compare RCFM fits to RAR and dark matter fits to the SPARC sample. The RAR approaches rotation curves from the MONDian paradigm of a changing law of inertia, but has also been shown to be consistent with dark matter fits on a sample of simulated galaxies (Keller & Wadsley, 2017).
@@ 2.4 New rotation curve fitting model (RCFM)
The physical paradigm for the new rotation curve fitting model (RCFM) is: the luminous mass is the only mass, the velocity due to the baryons is the only velocity, and excesses in Doppler shifted spectra represented by the “flat” rotation velocities are due to relative frame effects from the Milky Way galaxy. For clarity in what follows, we now drop explicit functional dependence on and set constants like Newton’s and the speed of light to 1. All terms can be assumed to be evaluated one-to-one in radius between the galaxy being observed and the Milky Way, with the exception of the model’s free parameter which is single valued for each galaxy fitted.
We construct a new heuristic rotation curve fitting formula by replacing the dark matter contribution to the rotation curve formula in Eq. 6 with the convolution of two Lorentz-type maps and which represent the frame-dependent effects due to the Milky Way;
This prescription is an implicit assumption that contributions to Doppler shifted spectra from translation ( in Eq. 5) and curvature () are roughly separable (Jackson, 1999) as represented by Eq. 7.
@@
In the frame field formalism “when going from one local inertial frame at a given point to another at the same point, the fields transform with respect to a Lorentz transformation” (Jetzer, 2017). Our goal then, is to use Lorentz-type transformations to map target galaxies onto the Milky Way, one-to-one in radius. To do this, we use the weak-field Schwarzschild time metric coefficients to characterize the frame fields, and hence the effects on photon frequency. The weak-field timelike Schwarzschild field frames are defined by
Terms in are the Newtonian gravitational potential (Wald, 1984). Classically, the of galaxies are integrated from the large limit of the data with a boundary value of , into a negative maxima at the small limit (Eq. 2). This is an implicit assumption of a globally flat embedding space, which ensures that the potential goes to zero at a distance of infinity from the mass distribution.
However, recent observations (Pomarède et al., 2020) have shown that on the relevant length scales of galaxies and groups of galaxies we have not recovered flat spacetimes. In fact, the external environments of galaxies and flowlines of groups of galaxies are exceedingly complex and diffusely populated with matter, demonstrating structure with sources and sinks. In addition, the value of the vacuum energy remains an outstanding physics problem (Zel’dovich, 1968).
In the absence of definite knowledge of the values of at the large limit of individual galaxies,
and because all RCFM terms are ratios of terms in for the emitter and receiver galaxies, we instead want to compare galaxies from a position where we know they have the same curvature, namely at their centers. To inform the integration of the gravitational potentials, we note Wolfgang Rindler’s statement that “the center of each galaxy provides a basic local standard of nonacceleration … so then can be treated like a local inertial frame relative to its own center”(Rindler, 2013).
From Rindler’s insight, we propose that to compare galaxies as inertial frames by Lorentz-type maps we must integrate from their centers out to the limit of the data.
In practice, this means that galaxy potentials are summed from at out to an essentially constant value
at the large limit of the data. This produces positive definite gravitational potentials,
which are then subtracted from in the Schwarzschild clock term .
Gravitational
potentials calculated in this way still obey Poisson’s equation Eq. 3 and the central force law Eq. 4, but allow for a non-flat embedding space with no loss of generality.
The RCFM mapping is accomplished in two steps. First, the curved 2-frame mapping maps the entire emitter galaxy onto the Milky Way one-to-one in radius,
for a rapidity defined by
where represents the emitter galaxy and the Milky Way. A cartoon of the transformation is represented in Fig. 2.
The second transformation takes this curved 2-frame (Eq. 12) into the flat field frames where observations are made and where the metric is Minkowski , as
for a rapidity angle defined by
for a convolution of the curved frame (Eq. 12) with the flat frame
Terms in are the Keplerian velocities from luminous mass as in Eq. 5. A cartoon of the transformation is represented in Fig. 3 .
That the Keplerian velocities are our best estimate of flatness is evidenced by the fact that dark matter is not required to reproduce the RC of our Solar System.
Previously, Rindler extended the transforms of special relativity to accelerated frames on a flat background (Boyer, 2011). Viewed as Rindler’s accelerated coordinates, and respectively can be seen to be timelike and spacelike transformations (Misner et al., 1973; Wald, 1984; Rindler, 2013).